OpenStudy (calculusxy):

Use the point-slope form equation to solve this problem: Passing through (3,5) and (3, 2)

OpenStudy (tommynaut):

The point-slope equation is \[y - y _{1} = m(x - x _{1})\] where \[(x_{1}, y_{1})\] is any point on the line, and m is the gradient. To solve this problem, use the gradient formula \[m = \frac{ y_{2} - y_{1} }{x_{2} - x_{1} }\] first to find m, then choose either point to plug into the point-slope equation.

OpenStudy (calculusxy):

I got m to be 0.

OpenStudy (tommynaut):

Are you sure you got 0 as the gradient? A negative denominator doesn't mean zero, it means... well, you can't divide by 0, so some people would say it means infinity. In this case though, a line with infinite slope would just be a vertical line, right?

OpenStudy (tommynaut):

A zero denominator, I mean. Sorry for that confusion.

OpenStudy (calculusxy):

So what am I supposed to put for that?

OpenStudy (tommynaut):

I was leading you down the wrong path :P When we have two point that lie on either a horizontal line or a vertical line, it becomes much easier to find the equation! No formulas needed. If you were to roughly plot the two points, what do you notice about the line?

OpenStudy (calculusxy):

They would be on the same x-axis

OpenStudy (calculusxy):

@Tommynaut

OpenStudy (calculusxy):

So would it be undefined slope?

OpenStudy (tommynaut):

Yes, they would be on the same vertical line. Remember, all horizontal lines are of the form y = b, where b is some number; and all vertical lines are of the form x = a, where a is some number. So we know our equation is going to look like x = a. The question is, what is a? What can we tell about all the x values on our line? And yes, the slope is undefined.

OpenStudy (tommynaut):

So, what did you get as your final answer?